OpenStudy (anonymous):

how do i go from 2x^2/sqrtx^2+1 + sqrt of x^2+1 to 4x2/sqrtX^2+1

OpenStudy (anonymous):

I think that 4is a 3, you get a common denominator and add, that's it

OpenStudy (anonymous):

yeah that what i get but the the book says the answer is 4x^2+2 /sqrtx^2+1

OpenStudy (anonymous):

Sometimes books are wrong. If all you doing is rewriting it that all that can be done to compact it

OpenStudy (anonymous):

yeah i know i also tried that wolfram site said the same thing

OpenStudy (anonymous):

thansk for the help rou i appreciate it

OpenStudy (anonymous):

thanks*

OpenStudy (anonymous):

\[\frac{2x^2}{\sqrt{x^2+1}} + \sqrt{x^2+1} \] \[= \frac{2x^2}{\sqrt{x^2+1}} + \frac{x^2 + 1}{\sqrt{x^2+1}}\] \[= \frac{3x^2+1}{\sqrt{x^2+1}} \] If the book says otherwise then either you've written the problem down wrong, or the book is wrong.